OpenStudy (anonymous):

explain how to find the equation for the quadratic function that has its vertex at (-1,2) and contains the point (0,-1)

ganeshie8 (ganeshie8):

you know the vertex form of quadratic equation

ganeshie8 (ganeshie8):

its a(x-h)^2 + k vertex = (h,k)

ganeshie8 (ganeshie8):

so with a vertex at (-1, 2) the equation becomes a(x-(-1))^2 + 2 = a(x+1)^2 + 2

ganeshie8 (ganeshie8):

you need to find "a"

ganeshie8 (ganeshie8):

you have some idea how to find "a" ? (check if any other info is given in the question other than vertex)

OpenStudy (anonymous):

thanks it does ask for nothing else

ganeshie8 (ganeshie8):

hmm the equation is incomplete you need to find "a" and plugin the value

OpenStudy (anonymous):

oh well the question is weird than